OpenStudy (anonymous):

Which is the equation of the given line in standard form? (Points : 1) –8x + 9y = 23 –8x + 9y = –23 –8x + 7y = 25 –9x + 8y = –23

OpenStudy (anonymous):

@mathmath333 @mathstudent55 @satellite73 @Hero

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

can u please help i don't have much time :(

OpenStudy (mathmath333):

(x1,y1)=(-4.-1) (x2,y2)=(1/2,3) y-y1=m(x-x_1) m=(y2-y1)/(x2-x1) plug in the values

OpenStudy (anonymous):

how do i wright it in standard form?

OpenStudy (mathmath333):

first plug in the values then rearrange